# Large-scale Generative AI Models Lack Visual Number Sense

### v1

@@ Abstract
Humans can readily judge the number of objects in a visual scene, even without counting, and such a skill has been documented in a variety of animal species and in babies prior to language development and formal schooling. Numerical judgments are error-free for small sets, while for larger collections responses become approximate, with variability increasing proportionally to the target number. This response pattern is observed for items of all kinds, despite variation in object features (such as color or shape), suggesting that our visual number sense relies on abstract representations of numerosity. Here, we investigated whether generative Artificial Intelligence (AI) models based on large-scale transformer architectures can reliably name the number of objects in simple visual stimuli or generate images containing a target number of items in the 1-10 range. Surprisingly, none of the foundation models considered performed in a human-like way: They all made striking errors even with small numbers, the response variability often did not increase in a systematic way, and the pattern of errors varied with object category. Our findings demonstrate that advanced AI systems still lack a basic ability that supports an intuitive understanding of numbers, which in humans is foundational for numeracy and mathematical development.
Keywords Foundation Models  $\cdot$ Machine Vision  $\cdot$ Deep Learning  $\cdot$ Numerical Cognition  $\cdot$ Numerosity
@@ Introduction
Artificial Intelligence (AI) is progressing rapidly, with deep learning models approaching or even surpassing human performance in a variety of domains, including perceptual judgements [1] and natural language processing [2]. In this article, we investigate whether advanced AI systems can judge the numerosity of visual sets, a core capability that we share with many animal species [3]. Even infants are sensitive to numerosity [4] and toddlers can generate sets containing a target number of items [5], suggesting that a preverbal understanding of numerical quantities develops well before formal education. Although it is a high-level feature derived from global image properties, numerosity is spontaneously extracted by our visual system [6]. Small numerosities in the “subitizing” range (up to 4) are perceived in an exact manner, while the numerosity of larger sets is approximately estimated when counting is precluded [7]. In the latter case, responses follow Weber’s law, so that variability increases proportionally to the mean estimate [3]. Another key signature of number sense is its abstract nature: numerosity is encoded independently from object category, location or presentation modality [4].
@@
Here we systematically probed the visual number sense of several state-of-the-art “foundation models”, which are large-scale generative architectures trained on huge data sets that can be readily used to solve a wide range of downstream tasks [8]. In the image-to-text domain, modern Visual Question Answering (VQA) systems can provide written answers to non-trivial questions about the content of an image or accurate descriptions of complex visual scenes; here we tested the capability of judging visual numerosity of the Vision-and-Language Transformer (ViLT) [9] and the Bootstrapping Language-Image Pre-training (BLIP-2) model [10]. In the text-to-image domain, generative models can produce high-quality visual content following detailed user prompts provided in natural language; here we tested the numerosity production skills of two of the most popular generative models for images: DALL-E (version 2) [11] and Stable Diffusion (version 2.1) [12].
@@
In line with recent proposals [13], we exploited two behavioral tasks used in cognitive psychology to evaluate number sense: numerosity naming [7], which requires establishing how many items are present in a given image, and numerosity production [14, 5], which requires generating a target number of items. We characterized the distribution of responses produced by the different AI models in the range 1-10 using a variety of object categories. Perfect accuracy across the entire numerical range would suggest the emergence of counting skills, while error-free responses with only small numbers would either indicate subitizing capabilities, or that counting is only partially developed as in children who do not fully master the counting principles [15, 16]. Error-prone responses centered on the target number would instead suggest that the AI model relies on approximate estimation mechanisms, which may or may not follow Weber’s law.
@@ Results
We probed the numerosity naming skills of two state-of-the-art VQA architectures: the Vision-and-Language Transformer (ViLT) [9] and the Bootstrapping Language-Image Pre-training (BLIP-2) model [10]. These systems have remarkable visual reasoning abilities and can answer nontrivial questions related to image content (e.g., What does the image represent? What are the feelings of the people in the scene and why?). We asked the models how many objects were present in a set of images containing up to 10 objects (see Section Materials and Methods for methodological details). Each stimulus included only items of the same kind, sampled from 5 possible classes: realistic pictures of common categories (apples, people, butterflies), colored dots, and “fast cards” depicting regularly placed clip-arts similar to those used to test children [15]. Examples are shown in Fig. 1 (top row).
Figure 1: Examples of images provided as input or produced by the models, ordered according to target number (reported at the bottom). The first row shows examples of stimuli used in the numerosity naming task: the first four images are taken from the “apples”, “people”, “butterflies” and “dots” categories, while the remaining images are sampled from the “fast cards” category. Subsequent rows show examples of images generated by DALL-E and Stable Diffusion in the numerosity production task, showcasing both correct and wrong generations. We sampled two images for each target category (apples, people, butterflies, dots). In specific cases, DALL-E never makes mistakes when there are a few objects to be produced (up to 3 objects for the “apples” category, exactly 3 objects for the “people” category, and exactly 1 object for the “butterflies” category). Another interesting phenomenon is the generation of images containing a wrong number of dots, which are nevertheless arranged according to the target digit shape (e.g., 8 in the figure).
@@
In general, response accuracy was extremely low for both models (ViLT: 28.0%, BLIP-2: 29.6%), thereby showing that they cannot count. In sharp contrast with humans, both models returned wrong answers within the subitizing range (1-4) and even for images with one or two objects. According to standard criteria used in developmental studies (see Methods), these models could be considered at most “One”-knowers [15]. Confusion matrices (CMs) reported in Fig. 2 clearly show the presence of anchoring effects, leading the models to choose prototypical responses (e.g., 4, 6) probably corresponding to common spatial layouts that might be over-represented in the training corpora. The pattern of responses also drastically varied between categories (minimum correlation between CMs for ViLT: 0.04, BLIP-2: 0.27), suggesting that the emergent numerical representations are not abstract.
@@
We also probed the numerosity production skills of two of the most popular generative models for images: DALL-E (version 2) [11] and Stable Diffusion (version 2.1) [12]. These foundation models have proven capable of generating high-quality images following a textual description, also taking into account stylistic instructions, fine-grained details, and relational features (e.g., A photo of an astronaut riding a horse in photorealistic style). We asked the models to generate images with a target number of objects, in analogy with numerosity production tasks used in animal and human studies [14, 5]. The target objects belonged to the same classes used for the naming task, except for the “fast cards” category, which could be underrepresented in the corpora used to train these models. Examples of generated images are shown in the middle and bottom rows of Fig. 1.
@@
Overall, DALL-E tended to generate images with white background and more definite objects, while Stable Diffusion adopted more artistic generation styles, producing either paintings, clip-arts, or realistic content. Similarly to the naming task, mean accuracy was low (DALL-E: 38.7%, Stable Diff: 33.3%), though DALL-E in some cases exhibited accurate response in the small number range (for the apples category, responses were error-free up to $n$ = 3). According to standard criteria, DALL-E would be considered a “Three”-knower, while Stable Diffusion would be considered a “Two”-knower [15]. Compared to the naming task, the response patterns were more homogeneous across categories (minimum correlation between CMs for DALL-E: 0.76, Stable Diff: 0.81).
Figure 2: Confusion matrices showing the distribution of models’ responses for different object categories. The x-axis represents the target number, while the y-axis represents the corresponding model responses. Response frequency is encoded using a grey scale (white = 0%, black = 100%). The first two rows refer to the numerosity naming task, which contains the additional “fast cards” category. The remaining rows refer to the numerosity production task, and contain two additional panels depicting the response distributions observed in behavioral studies with human adults [7] and children [16] across similar numerical ranges (note that for the adult data the maximum number of items to be estimated was eight). For the children data, the graph shows the response frequency of “Three”-knower children (larger circles indicate higher frequency), along with the estimated response likelihood (gray shading).
Adherence to Weber’s law
@@
In humans and other animal species, numerosity estimation data collected through naming and production tasks approximately follow the law of scalar variability, as indicated by a constant coefficient of variation (standard error of the responses/mean response) [17, 18] (but see [19] for discussion of possible exceptions). To investigate whether models’ responses followed a scalar variability pattern, we used a well-established methodology implementing a log–log regression analysis [20]. If scalar variability holds, as the number $n$ increases the logs of the standard error $SE(n)$ and of the mean estimates $\mu(n)$ should form a straight line, with a slope $\beta=1$ [$log(SE(n))=\alpha+\beta log(\mu(n))$]. A slope that is significantly different from 1 would indicate that the relation between the standard errors and the means is a more complex power function, while a poor fit would more generally indicate that the response variability cannot be explained by linear or power trends.
@@
We carried out a separate analysis for each model and for each object category. If a model exhibited error-free responses for small numbers, we excluded those numbers from the analysis. In almost half of the cases (8 out of 18), the regression analysis resulted in a poor fit ($p>0.05$), indicating that the distribution of responses did not systematically change as a function of numerosity, in sharp contrast with human and animal data. In the remaining cases the regression analysis was significant (all $p<0.05$; $r^{2}=0.64-0.96$), therefore we further tested whether the regression slopes were significantly different from 1. Results were mixed, and depended both on the model type and the object category.
@@
In the numerosity naming task, the distribution of models’ responses did not follow the pattern observed in human estimation [7]. For ViLT, we observed scalar variability for the people (slope = 1.11, p > 0.05) and butterflies (slope = 1.14, p > 0.05) categories, while we observed power variability for the fast card category (slope = 1.37, p = 0.003). For BLIP2, the regression analysis never produced a good fit. In the numerosity production task the errors were more regularly distributed around the target number (also see Fig. 2). For DALL-E, we observed scalar variability for the dots category (slope = 0.82, p > 0.05), while we observed power variability for the apples (slope = 1.53, p = 0.024) and butterflies (slope = 1.49, p = 0.036) categories. For Stable Diffusion, we observed scalar variability for the apples (slope = 1.07, p > 0.05) and butterflies (slope = 1.15, p > 0.05) categories, while we observed power variability for the people (slope = 0.70, p = 0.030) and dots (slope = 0.60, p = 0.016) categories.
@@
Overall, scalar variability was thus observed only in five specific conditions. In other five conditions the variability increased according to a power law, while in the remaining eight cases the response variability did not have a systematic trend.
@@ Discussion
This work demonstrates that foundational AI models cannot reliably judge the number of objects in a visual scene, both in image-to-text and text-to-image tasks. Such a striking deficit is observed even for small sets, supporting a knower-level comparable to that of 3-year-old children [15, 16]. This might explain anecdotal evidence showing the failure of generative AI to synthesize realistic images featuring multiple instances of a body part, such as the correct number of fingers in a hand. Surprisingly, and in contrast with smaller-scale deep learning models [21, 22, 23], even approximate estimation was not robust and often failed to match the psychophysics of human numerosity perception. We believe that substantial progress in architecture design and training procedures is warranted to create AI systems that understand visual numerosity [24]. Grounding numeracy development in the visual number sense, as in the case of humans [3], might also be the key to enabling AI systems to acquire and fully master numerical and mathematical knowledge [25].
@@
In the naming task, for each object class and target number we probed the models using 50 high-resolution ($1024\times 1024$ pixels) images created by randomly placing items of variable size on a uniform white background, with no overlap. Fast card stimuli were created using clip-arts of common objects (apples, bells, butterflies, candies, cars, fish, flowers, planes, stars) drawn in different colors (black, blue, green, orange, red). Responses without numerical content or with vague quantification (e.g., “a few”) were automatically discarded, but we ensured that at least 20 acceptable answers were produced for each class/number combination. In the production task, for each condition we probed the models to generate 50 high-resolution images, which were then manually labeled by one of the authors (K.H.) and individually double checked by another author. Images with an ambiguous number of objects or with ill-formed content were discarded (see Section Analysis of model responses for details). We ensured that at least 20 acceptable images were produced for each test condition.
@@
We considered two different Visual Question Answering (VQA) models. We used the vilt-b32-mlm version of the Vision-and-Language Transformer (ViLT) model [9] available through Hugging Face. This architecture incorporates text embeddings into a Vision Transformer (ViT), allowing it to have a minimal design for Vision-and-Language Pre-training and thus speeding-up model training and inference phases. It has a total of 87.4 million parameters. We also used the blip2-flan-t5-xl version of the Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP-2) model [10], developed by Salesforce and available through Hugging Face. This architecture is considered an improved version of BLIP [26] and it is available in different versions, depending on the type and size of the backbone models. We explored all versions except for the t5-xxl model (due to GPU memory constraints) and found that the version using Flan-T5 as a language model yielded the best accuracy. The chosen model version has a total of 4.1 billion parameters.
We considered two different image generation models. We probed the DALL-E (version 2) model [11] using the API interface provided by OpenAI. DALL-E 2 is an improved version of the original text-to-image DALL-E model, featuring a total of 3.5 billion parameters. We also used the Stable Diffusion (version 2.1) model [12], developed by Stability AI and available through Hugging Face. Stable Diffusion v2.1 is a text-to-image latent diffusion model that combines an autoencoder with a diffusion model that is trained in the latent space of the autoencoder. It has approximately 500 million parameters.
@@
For the numerosity naming task, we implemented three different prompting methods. The first required to explicitly estimate the number of objects belonging to a specific category (i.e., How many apples / butterflies / people / dots / shapes are there in the picture?). The other two used the more general “objects” or “things” words to identify the items to count. For ViLT, the best performance was achieved with the generic “things” prompt, while for BLIP-2 the best performance was achieved with the category-specific prompts. To make sure that the models were prompted correctly, we carried out a control simulation related to a non-numerical task using the entire set of “apples” stimuli. We probed the VQA models with the following prompt: What does the image represent? and we considered as correct the following answers: apple(s) and fruit. All models almost always provided the correct answer across the entire set of stimuli (accuracy for ViLT : 96.8%, BLIP-2 : 100%), thus demonstrating a proper understanding of the image content and the prompt structure.

### v2

# Visual Enumeration is Challenging for Large-scale Generative AI

@@ Abstract
Humans can readily judge the number of objects in a visual scene, even without counting, and such a skill has been documented in many animal species and babies prior to language development and formal schooling. Numerical judgments are error-free for small sets, while for larger collections responses become approximate, with variability increasing proportionally to the target number. This response pattern is observed for items of all kinds, despite variation in object features (such as color or shape), suggesting that our visual number sense relies on abstract representations of numerosity. Here, we investigate whether large-scale generative Artificial Intelligence (AI) systems have a human-like number sense, which should allow them to reliably name the number of objects in simple visual stimuli or generate images containing a target number of items in the 1-10 range. Surprisingly, most of the foundation models considered have a poor number sense: They make striking errors even with small numbers, the response variability does not increase in a systematic way, and the pattern of errors depends on object category. Only the most recent proprietary systems exhibit signatures of a visual number sense. Our findings demonstrate that having an intuitive visual understanding of number remains challenging for foundation models, which in turn might be detrimental to the perceptual grounding of numeracy that in humans is crucial for mathematical learning.
Keywords Foundation Models  $\cdot$ Machine Vision  $\cdot$ Deep Learning  $\cdot$ BLIP  $\cdot$ ViLT  $\cdot$ GPT  $\cdot$ Gemini  $\cdot$ DALL-E  $\cdot$ Stable Diffusion  $\cdot$ Numerical Cognition  $\cdot$ Numerosity  $\cdot$ Counting
@@ Introduction
Artificial Intelligence (AI) is progressing rapidly, with deep learning models approaching or even surpassing human performance in a variety of domains, including perceptual judgements [1] and natural language processing [2]. In this article, we investigate whether advanced AI systems can judge the numerosity of visual sets, a core capability that humans share with many animal species [3]. Even infants are sensitive to numerosity [4] and toddlers can generate sets containing a target number of items [5], suggesting that a preverbal understanding of numerical quantities develops well before formal education. Small numerosities in the “subitizing” range (up to 4) are perceived in an exact manner, while the numerosity of larger sets is approximately estimated when counting is precluded [6]. In the latter case, responses follow Weber’s law, so that variability increases proportionally to the mean estimate [3]. Another key signature of number sense is its abstract nature: numerosity is encoded independently from object category, location or presentation modality [4]. Importantly, numerosity is spontaneously extracted by our visual system [7] and there is broad consensus that numerosity perception is foundational for subsequent learning of symbolic numbers as well as for the acquisition of higher-level mathematical competence [8, 3].
Machine vision researchers have engineered a variety of specialized systems for visual object counting, often tailored to specific categories such as penguins [9] or crowds [10]. The most popular framework consists on first running an object detector to segment the target items in a visual scene and then explicitly counting the resulting bounding boxes or object proposals [11, 12], often summing fractional counts estimated from different sections of the image [13]. However, in these approaches numerosity representations do not emerge within the model itself, because the encoding of number is delegated to an external, often hard-wired mechanism. A different perspective considers the possibility that numerosity representations might spontaneously emerge in deep neural networks as a high-order statistical feature of the sensory environment [14]. Indeed, it has been shown that a rudimentary visual number sense can emerge in small-scale generative models trained with the goal of reconstructing images with a varying number of objects [15, 16].
In this work, we investigate whether numerosity perception abilities spontaneously emerge in state-of-the-art generative AI systems. To this end, we systematically probe the visual number sense of several “foundation models”, which are large-scale generative architectures trained on huge data sets that have shown emergent abilities in a variety of domains [17] and can readily solve a wide range of downstream tasks [18]. Unlike the domain-specific architectures mentioned above, foundation models are domain-general systems that can be used out-of-the-box without the need of fine-tuning on numerical tasks. However, despite their flexibility and their remarkable performance in a variety of problems, it has been repeatedly shown that foundation models often fall short in tasks that require the manipulation of numerical information [19].
@@
We investigate key properties of visual number sense across a range of models of different sizes and complexities. In the image-to-text domain, we consider modern Visual Question Answering (VQA) systems that can provide written answers to non-trivial questions about the content of an image or accurate descriptions of complex visual scenes. In particular, we test the capability of judging visual numerosity of the Vision-and-Language Transformer (ViLT) [20] and the Bootstrapping Language-Image Pre-training (BLIP-2) model [21]. We also consider two recent proprietary models, GPT-4V [22] developed by OpenAI and Gemini [23] developed by Google, which are considered the most advanced multimodal AI systems currently available. In the text-to-image domain, we consider generative models that can produce high-quality visual content following detailed user prompts provided in natural language. We test the numerosity production skills of popular generative architectures for images: Stable Diffusion (version 2.1) [24] and DALL-E (version 2 and version 3) [25, 26].
@@
In line with the proposal of using methods from cognitive science to test foundation models [27], we exploit two behavioral tasks that are widely used to evaluate number sense in humans: numerosity naming [6], which requires establishing how many items are present in a given image, and numerosity production [28, 5], which requires generating a target number of items. We characterize the distribution of responses produced by the different AI models in the range 1-10 using a variety of object categories. Perfect accuracy across the entire numerical range would suggest the emergence of counting skills, while error-free responses with only small numbers would either indicate subitizing capabilities, or that counting is only partially developed as in children who do not fully master the counting principles [29, 30]. Error-prone responses centered on the target number would instead suggest that the AI model relies on approximate estimation mechanisms, which may or may not follow Weber’s law.
@@ Results
We probed the numerosity naming skills of the Vision-and-Language Transformer (ViLT) [20], the Bootstrapping Language-Image Pre-training (BLIP-2) model [21], the latest multimodal version of the Generative Pre-trained Transformer (GPT-4V) [22] and the recently introduced multimodal Gemini model [23]. These systems have remarkable visual reasoning abilities and can answer non-trivial questions related to image content (e.g., What does the image represent? What are the feelings of the people in the scene and why?). We asked the models how many objects were present in a set of images containing up to 10 objects (see Materials and Methods for methodological details). Each stimulus included only items of the same kind, sampled from 5 possible classes: realistic pictures of common categories (apples, people, butterflies), colored dots, and “fast cards” depicting regularly placed clip-arts similar to those used to test children [29]. Examples of visual stimuli are shown at the top of Fig. 1.
Figure 1: Confusion matrices for the numerosity naming task. Each panel shows the distribution of models’ responses across different object categories: apples, people, butterflies, dots and fast cards (a few sample stimuli are shown below each category name). The x-axis represents the target number, while the y-axis represents the corresponding model responses. Response frequency is encoded using a perceptually uniform colormap (blue = 0%, yellow = 100%).
@@
In general, the response accuracy was extremely low for both ViLT and BLIP models (ViLT: 28.0%, BLIP-2: 29.6%), indicating that they cannot count. Moreover, in sharp contrast with humans, ViLT and BLIP models also returned wrong answers within the subitizing range (1-4) and even for images with one or two objects. According to standard criteria used in human developmental studies (see Methods), these models could be considered at most “One”-knowers, that is, they exhibit reliable enumeration only for a single object, as typical of children younger than three years of age [29]. Confusion matrices (CMs) reported in Fig. 1 clearly show the presence of anchoring effects, leading the models to choose stereotyped responses (e.g., 4, 6) probably corresponding to common spatial layouts that might be over-represented in their training corpora. The pattern of responses also drastically varied between categories (minimum correlation between CMs for ViLT: 0.04, BLIP-2: 0.27), suggesting that they fail to abstract numerical information.
The responses of GPT-4V and Gemini were much more accurate (73.9% and 70.6%, respectively), suggesting that these multimodal models might possess rudimentary enumeration skills. Responses were often error-free in the subitizing range, and confusion matrices were fairly consistent between categories (the minimum correlation was 0.87 for GPT-4V and 0.70 for Gemini). Considering the distribution of response errors, both GPT-4V and Gemini would be characterized as “Six”-knowers. Interestingly, such a pattern is never observed in human numeracy development because children typically transition from a “Four”-knower to a “Cardinal-principle”-knower level, which implies full mastery of the counting principles with no strict upper limit [29].
@@
We probed the numerosity production skills of three popular generative AI systems for images: Stable Diffusion (version 2.1) [24] and DALL-E (both versions 2 and 3) [25, 26]. These foundation models have proven capable of generating high-quality images following a textual description, also taking into account stylistic instructions, fine-grained details, and relational features (e.g., A photo of an astronaut riding a horse in photorealistic style). We asked the models to generate images with a target number of objects, in analogy with numerosity production tasks used in animal and human studies [28, 5]. Target objects belonged to the same classes used for the naming task, except for the “fast cards” category, which could be underrepresented in the corpora used to train these models. Examples of generated images are shown in Fig. 2, while confusion matrices are shown in Fig. 3.
Figure 2: Examples of images generated by DALL-E 2 and Stable Diffusion in the numerosity production task, showcasing both correct and wrong generations (the target number is indicated at the bottom). We report two images for each target category: apples, people, butterflies, and dots. For some categories, DALL-E 2 never made mistakes when there were a few objects to produce. For the dots category, in a few cases DALL-E 2 generated images containing a wrong number of dots, which were nevertheless arranged according to the target digit shape (e.g., 8 in the figure).
Figure 3: Confusion matrices for the numerosity production task. The x-axis represents the target number, while the y-axis represents the corresponding model responses. Response frequency is encoded using a perceptually uniform colormap (blue = 0%, yellow = 100%).
@@
Overall, DALL-E 2 tended to generate images with a white background and more definite objects, while Stable Diffusion adopted more artistic generation styles, producing either paintings, clip-arts, or realistic content. DALL-E 3 produced more detailed images with respect to its predecessor model. The mean response accuracy was fairly low for all models (Stable Diff: 33.3%, DALL-E 2: 38.7%, DALL-E 3: 47.7%). Interestingly, DALL-E 2 was the only model capable of exhibiting error-free responses in the small number range, but only in specific cases (up to 3 objects for the “apples” category, exactly 3 objects for the “people” category, and exactly 1 object for the “butterflies” category). Despite its higher average accuracy, DALL-E 3 always produced a few errors even in the subitizing range. According to criteria used in human developmental studies, Stable Diffusion would be considered a “Two”-knower, DALL-E 2 a “Three”-knower, and DALL-E 3 a “Four”-knower. Compared to the naming task, the response patterns were more homogeneous across categories (minimum correlation between CMs for Stable Diff: 0.81, DALL-E 2: 0.76, DALL-E 3: 0.93).
Analysis of response distribution
@@
In humans and other animal species, numerosity tasks yield a distribution of error responses that is not random, but varies systematically according to Weber’s law [31]. In particular, the standard error of the response increases proportionally to the mean, indicating scalar variability, which can be indexed by a constant coefficient of variation (standard error of the responses/mean response) [32, 33]. To investigate whether models’ responses followed a scalar variability pattern, we used a well-established methodology implementing a log–log regression analysis [34]. If scalar variability holds, as the number $n$ increases the logs of the standard error $SE(n)$ and of the mean estimates $\mu(n)$ should form a straight line, with a slope $\beta=1$ [$log(SE(n))=\alpha+\beta log(\mu(n))$]. A slope that is significantly different from 1 would indicate that the relation between the standard errors and the means is a more complex power function, while a poor fit would more generally indicate that the response variability cannot be explained by linear or power trends.
@@
We carried out a separate analysis for each model and for each object category. If a model exhibited error-free responses for small numbers, we excluded those numbers from the analysis. In many cases the regression analysis resulted in a poor fit ($p>0.05$), indicating that the distribution of responses across numerosities did not match the trend observed in human and animal experiments. As detailed below, in the remaining cases the regression analysis was significant (all $p<0.05$; $r^{2}=0.64-0.96$), so we further tested whether the regression slopes were significantly different from 1. Results were mixed, and depended both on the model type and the object category.
@@
In the numerosity naming task, for ViLT and BLIP the distribution of models’ responses did not generally follow the pattern observed in human studies. For ViLT, we observed scalar variability only for the people (slope = 1.11, $p>0.05$) and butterflies (slope = 1.14, $p>0.05$) categories, while we observed power variability for the fast cards category (slope = 1.37, $p=0.003$). For BLIP2, the regression analysis never produced a good fit. For GPT-4V, we observed scalar variability for the people (slope = 0.98, $p>0.05$), butterflies (slope = 1.07, $p>0.05$) and fast cards (slope = 1.23, $p>0.05$) categories, while we observed power variability for the dots category (slope = 0.73, $p=0.008$). For Gemini, the coefficient of variation was not stable across numerosities, resulting in a poor fit for almost all categories (p > 0.05). Only for the people category we observed a systematic power variability trend (slope = 0.62, $p=0.033$).
In the numerosity production task the errors were more regularly distributed around the target number. For Stable Diffusion, we observed scalar variability for the apples (slope = 1.07, $p>0.05$) and butterflies (slope = 1.15, $p>0.05$) categories, while we observed power variability for the people (slope = 0.70, $p=0.030$) and dots (slope = 0.60, $p=0.016$) categories. For DALL-E 2, we observed scalar variability for the dots category (slope = 0.82, $p>0.05$), while we observed power variability for the apples (slope = 1.53, $p=0.024$) and butterflies (slope = 1.49, $p=0.036$) categories. DALL-E 3 was the model most closely adhering to Weber’s law, since we observed scalar variability across all categories: apples (slope = 0.98, $p>0.05$), people (slope = 1.10, $p>0.05$), butterflies (slope = 0.85, $p>0.05$) and dots (slope = 0.98, $p>0.05$).
@@
Overall, scalar variability was thus mostly observed only in the most recent OpenAI models, GPT-4V and DALL-E 3. In the other models the response variability often increased according to a power law or did not have a systematic trend.
@@ Discussion
This work demonstrates that large-scale foundation models cannot yet reliably enumerate the number of objects in a visual scene, both in image-to-text and text-to-image tasks. Such a striking deficit is observed even for sets containing just a few items, suggesting a number knower-level that is at best comparable to that of preschool children that do not fully master the counting principles [29, 30]. This might explain anecdotal evidence showing the failure of generative AI to synthesize realistic images featuring multiple instances of a body part, such as the correct number of fingers in a hand.
Surprisingly, even approximate number estimation failed to match the psychophysics of human numerosity perception. Only the most recent proprietary models, GPT-4V and DALL-E 3 from OpenAI, exhibited sparks of human-like number sense: their responses were often error-free for small numbers, suggesting some subitizing capabilities, and sometimes followed scalar variability for larger numbers, in accordance to Weber’s law. Also Gemini, the most recent multimodal model from Google, achieved a fairly accurate performance in the numerosity naming task, with error-free responses in the subitizing range. However, the response variability for larger numbers in this case did not adhere to Weber’s law.
The fact that the concept of visual numerosity remains elusive for large-scale foundation models is particularly striking when considering that numerosity in humans is susceptible to adaptation effects [35], which is the hallmark of a primary perceptual property (just as orientation or color). Accordingly, neuronal populations encoding numerosity have been found in multiple cortical regions in human neuroimaging experiments [36, 37, 38] and in neurophysiological studies with animals [39]. Moreover, computational modeling studies have shown that sensitivity to numerosity can emerge in small-scale deep learning models trained to generate synthetic images of object sets [15, 16, 40]. Diffusion models, such as Stable Diffusion and DALL-E, are trained with a similar objective on huge and heterogeneous image datasets that most likely contain substantial variability in numerosity. However, the empirical distribution of visual numerosities in natural image datasets is approximately captured by Zipf’s law (that is, frequency is inversely proportional to number [41]), therefore, it might be possible that oversampling of small numerosities in the training datasets of foundation models has detrimental effects on their emergent representational space. This still does not explain the variability across object categories observed in our analyses, which suggests that the representation of numerosity is not fully disentangled from other image properties. An alternative explanation for the poor enumeration skills of foundation models lies in the mapping between perceptual numerosity representations, encoded in image embeddings, and number symbols (number words or Arabic digits) encoded in text embeddings. In children, establishing such a bidirectional mapping is a sophisticated developmental process, which takes many years and requires explicit instruction [42].
In this respect, it would be valuable to analyze and compare the embeddings emerging in different models to gain a deeper understanding of the computational principles that could enable the development of stronger enumeration skills. Unfortunately, the closed-source nature of proprietary systems does not allow us to draw strong conclusions about the nature of their putative visual number sense. The fact that performance was never perfect even for the most advanced models suggests that numerosity estimation was not engineered in the system, but we cannot exclude the possibility that some counting mechanisms were partially introduced as extra processing layers during prompt elaboration. Furthermore, the most sophisticated AI systems can also exploit the self-generation of code snippets to fulfill a user request, as in the case of mathematical problem solving [43]. In other words, an advanced AI system could in principle use external tools (e.g., based on object detection and a symbolic counting algorithm) to carry out visual enumeration without any understanding of numerosity and counting principles.
Whether numerosity representations could spontaneously develop as an emergent ability [17] in foundation models thus remains an important open question, which would require “opening the box” and inspecting the model’s inner functioning in a way that is not currently possible with proprietary models. This issue highlights the dangers of using proprietary models in academic research [44], and calls for further research efforts to develop open source foundation models with a focus on basic perceptual abilities, such as those underlying our visual number sense.
In conclusion, we believe that substantial progress in architecture design and training procedures is warranted to create AI systems that truly understand visual numerosity [45]. Grounding numeracy development in the visual number sense, as in the case of humans [3], might also be the key to enabling AI systems to acquire and fully master numerical and mathematical knowledge without resorting to highly specialized hybrid architectures [46].
@@
In the naming task, for each object class and target number we probed the models using 50 high-resolution ($1024\times 1024$ pixels) images created by randomly placing items of variable size on a uniform white background, with no overlap. Fast card stimuli were created using clip-arts of common objects (apples, bells, butterflies, candies, cars, fish, flowers, planes, stars) drawn in different colors (black, blue, green, orange, red). Responses without numerical content or with vague quantification (e.g., “a few”) were automatically discarded, but we ensured that at least 20 acceptable answers were produced for each class/number combination. In the production task, for each condition we probed the models to generate 50 high-resolution images, which were then manually labeled by one of the authors (K.H.) and individually double checked by another author. Images with an ambiguous number of objects or with ill-formed content were discarded (see Materials and Methods for details). We ensured that at least 20 acceptable images were produced for each test condition.
@@
For the numerosity naming task, we considered two different Visual Question Answering (VQA) models. We used the vilt-b32-mlm version of the Vision-and-Language Transformer (ViLT) model [20] available through Hugging Face. This architecture incorporates text embeddings into a Vision Transformer (ViT), allowing it to have a minimal design for Vision-and-Language Pre-training and thus speeding-up model training and inference phases. It has a total of 87.4 million parameters. We also used the blip2-flan-t5-xl version of the Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP-2) model [21], developed by Salesforce and available through Hugging Face. This architecture is considered an improved version of BLIP [47] and it is available in different versions, depending on the type and size of the backbone models. We explored all versions except for the t5-xxl model (due to GPU memory constraints) and found that the version using Flan-T5 as a language model yielded the best accuracy. The chosen model version has a total of 4.1 billion parameters.
We also considered the two most advanced multimodal models currently available, GPT-4V [22] and Gemini Pro [23], which can be readily used in VQA settings. These models are regarded among the most powerful generalist AI systems to date, thanks to their unprecedented ability to understand and process an arbitrary mix of input images and texts. Technical details regarding the underlying architecture and inner working of these models (including engineered modules that might be used to solve specific tasks) have not been revealed; it has been speculated that both these models might have more than one trillion of parameters [^1] https://the-decoder.com/gpt-4-architecture-datasets-costs-and-more-leaked.
For the numerosity production task, we considered two different image generation architectures. One is represented by the Stable Diffusion (version 2.1) model [24], developed by Stability AI and available through Hugging Face. Stable Diffusion v2.1 is a text-to-image latent diffusion model that combines an autoencoder with a diffusion model that is trained in the latent space of the autoencoder. It has approximately 500 million parameters. We also tested models from the DALL-E family (version 2 and version 3) [25, 26] using the API interface provided by OpenAI. DALL-E 2 is an improved version of the original text-to-image DALL-E model, featuring a total of 3.5 billion parameters. DALL-E 3 is the latest version, which was trained using highly descriptive, synthetic captions for the training images. The number of model parameters is currently unknown.
@@
For the numerosity naming task, we implemented three different prompting methods. The first required to explicitly estimate the number of objects belonging to a specific category (i.e., How many apples / butterflies / people / dots / shapes are there in the picture?). The other two used the more general “objects” or “things” words to identify the items to count. For ViLT, GPT-4V and Gemini, the best performance was achieved with the generic “things” prompt, while for BLIP-2 the best performance was achieved with the category-specific prompts.
To make sure that the models were prompted correctly, we carried out a control simulation related to a non-numerical task using the entire set of “apples” stimuli. We probed the VQA models with the following prompt: What does the image represent? and we considered as correct the following answers: apple(s) and fruit. All models almost always provided the correct answer across the entire set of stimuli (accuracy for ViLT : 96.8%, BLIP-2 : 100%, GPT-4V : 100%, Gemini: 100%), thus demonstrating a proper understanding of the image content and the prompt structure.